OpenStudy (eric_d):

sketch graph of rational function, showing asymtotes for following function: y= x^2/ 1+x vertical asymtote is -1 can anyone show me how to find the slant asymtote?

ganeshie8 (ganeshie8):

divide

ganeshie8 (ganeshie8):

\(\large \dfrac{x^2}{x+1} = \color{red}{x-1} + \dfrac{1}{x+1}\)

ganeshie8 (ganeshie8):

slant asymptote is : \( y = \color{red}{x-1} \)

OpenStudy (eric_d):

how do you get x-1

ganeshie8 (ganeshie8):

knw long division ?

OpenStudy (eric_d):

don't mind showing me

OpenStudy (eric_d):

Nvm, I know...! Thanks Ganes

ganeshie8 (ganeshie8):

np :) i can walk u thru the long division if u want...

ganeshie8 (ganeshie8):

u may graph ur rational function here to get a feel of how it looks, and it may help u sketch manually : https://www.desmos.com/calculator

OpenStudy (eric_d):

Thnks

ganeshie8 (ganeshie8):